Find the compound amount and the amount of interest earned by the following deposit. ​$3.000 at 6.59​% compounded continuously for 9 years.

What is the compound​ amount?

What is the interest earned?
Answer #1

Future Value = Present Value * e^(Rate per year*no. of years)

= 3000*e^(.0659*9)

= 3000*e^0.5931

= 3000*1.80958945661

= 5428.77

interest earned = Future Value - Present Value

= 5428.77-3000

= 2428.77